# Adaptive DRL for IRS Mirror Orientation in Dynamic OWC Networks

### v1

@@ V-B Simulation Results
Fig. 5 shows the bit error rate (BER) versus the SNR when IRS$=10\times 10$ are deployed, and different user velocities are considered. It can be seen that the BER increases with the speed of the user. For instance, when the SNR is 15 dB, the BER increases from $10^{-4}$ to $10^{-1}$. It is worth mentioning that as the user moves faster, it becomes more challenging for our algorithm to precisely track its location and steer the mirrors towards it. However, as the SNR increases, the proposed algorithm achieves a BER below $10^{-3}$.
@@ VI Conclusion
This paper explored the application of mirror-based IRS to improve signal coverage and mitigate LoS blockages in dynamic indoor VLC systems. We addressed the challenge of optimizing IRS mirror orientation by formulating the problem as a MDP and applying the DDPG algorithm, a powerful DRL technique. The proposed DDPG-based solution enables real-time adaptability to user mobility and environmental blockages. Simulation results validate the effectiveness of our approach, showing that it consistently outperforms conventional DRL methods and yields significant gains in sum rate over fixed-orientation IRS configurations.
Acknowledgment
This work has been supported by the Engineering and Physical Science Research Council (EPSRC), by the INTERNET project under Grant EP/H040536/1, by the STAR project under Grant EP/K016873/1, by the TOWS project under Grant EP/S016570/1, and by the TITAN project under Grant EP/X04047X/2. All data are provided in full in the results section of this paper.

### v2

@@ V-B Simulation Results
Fig. 4 shows the sum rate versus the optical transmitted power in different blockage scenarios. Our DRL algorithm is compared to a fixed orientation IRS as a benchmark scheme, where each IRS mirror points to a random direction. In other words, the IRS orientation in the benchmark scheme remains static throughout the serving time, regardless of user distributions and demands. It can be seen that the proposed DRL algorithm achieves up to 48% higher sum rates compared to the benchmark scheme when two blockages and moderate mobility (i.e., user velocity $\approx$ 1 m/s) are considered. This is due to the ability of the DRL algorithm to steer the mirrors every time the users move and/or encounter blockages. For instance, when the transmitted optical power is equal to 2 W, the proposed algorithm achieves almost 19.2 Mbps compared to 10 Mbps achieved by the benchmark scheme in the scenario of two blockages.
[…]
@@ VI Conclusion
This paper explored the application of mirror-based IRS to improve signal coverage and mitigate LoS blockages in dynamic indoor VLC systems. We addressed the challenge of optimizing IRS mirror orientation by formulating the problem as a MDP and applying the DRL algorithm, a powerful DRL technique. The proposed DRL-based solution enables real-time adaptability to user mobility and environmental blockages. Simulation results validate the effectiveness of our approach, showing that it consistently outperforms conventional DRL methods and yields significant gains in sum rate over random-orientation IRS configurations.